# GAI_arXiv.bib

### v2

@@ 3 Is retaining blades proposed, and is enclosing them inside the main body suggested? (1 point)
4 If no direct mention of a motor is made, is retaining blades proposed, or is a device as described in criterion 1 conventionally paired with a motor proposed? (1 point) Additionally, is the use of household electricity as the motor’s power source mentioned? (1 point)
5 Is retaining the motor proposed, and is it identified as playing a role in generating driving fluid or negative pressure? (1 point)
@@ 7 Is the introduction of a nozzle structure proposed? (1 point)
8 Other than the introduction of a nozzle structure, are any proposals made to optimize internal channels for improving the efficiency or noise reduction of the driving fluid? (Up to 2 points, depending on the number and specificity of the proposals)
@@ 9 Are relevant principles of fluid dynamics explicitly explained? (1 point)
To evaluate the performance of the proposed framework, the experimental results are compared across eight models shown in Table II. The eight models are differentiated based on three perspectives: (1) the number of agents, (2) the presence of internal states, and (3) the heterogeneity of intrinsic motivations. The first perspective examines the impact of increasing the number of agents participating in the dialogue. When multiple agents are involved, the evaluation is conducted based on the majority opinions. The second perspective assesses the role of the internal state module’s functionality. The third perspective evaluates the effect of whether the weights assigned to the three intrinsic motivations (novelty, importance, and consensus) are homogeneous or heterogeneous among agents. In homogeneous models (Models D and G), all agents equally consider the three motivations. In heterogeneous models (Models E and H), agents are configured with different emphases on intrinsic motivations: some prioritize novelty while de-emphasizing consensus, others prioritize consensus while de-emphasizing novelty, and some consider all three motivations equally.
Model Number of agents Internal state intrinsic motivation
A: 1 no-internal agent 1 No n.a.
B: 1 with-internal agent 1 Yes n.a.
C: 3 no-internal agents 3 No n.a.
D: 3 with-internal homogeneous agents 3 Yes homogeneous
E: 3 with-internal heterogeneous agents 3 Yes heterogeneous
F: 5 no-internal agents 5 No n.a.
G: 5 with-internal homogeneous agents 5 Yes homogeneous
H: 5 with-internal heterogeneous agents 5 Yes heterogeneous
TABLE II: Eight models for experiment
@@
The experiment is conducted 10 times for each of the eight models, using a common set of 10 seeds. The language model used to power the framework is OpenAI’s GPT-4o. To ensure uniform experimental conditions, word limits are imposed on the agents’ statements in each round. When expressing their own opinions, agents are limited to 170 words, while statements regarding the opinions of other agents are restricted to 50 words. In this preliminary experiment, a flat organizational structure was adopted, where all agents could communicate with all other agents. Other organizational structures, such as hierarchical models, are left for future research. In the evaluation, the ratings are done by carefully examining each agent’s statement on possible solutions, one by one, through human assessment.
III-D Results
@@
As shown in Table III, in all cases with 1, 3, and 5 agents, the models with internal states outperformed those without on average. This indicates that, through refinement by the internal state module, more coherent and comprehensive ideas are constructed and shared among agents. For example, comparing the two models with one agent, in the model without an internal state (Model A), the proposal of retaining blades or incorporating an appropriate blade-driven device as a mechanism to generate a driving fluid or negative pressure (Criteria 1 or 2) was made only 3 times out of 10. In contrast, in the model with an internal state (Model B), such mechanisms were proposed 9 times out of 10. This indicates, in the former model, the agent’s reflection did not address how to deal with the external driving fluid required in incorporating ejector-like technology into household fan, whereas in the latter model, the concept of external driving fluid was successfully reframed within the context of household fans as the intake of surrounding air by the fan blades. Furthermore, in the model without an internal state, aside from the introduction of nozzle structures, the proposal for optimizing internal channels to improve the efficiency and reduce the noise of the driving fluid (Criterion 8) was made only once out of 10 times. In contrast, in the model with an internal state, such proposals were made 7 times out of 10, ranging from the redesign of the fan housing to the introduction of guide vanes and the use of Computational Fluid Dynamics (CFD) simulations. Below is an example of part of an agent’s statement in the single-agent model with an internal state (Model B).
@@
To address challenges in household fans, we could integrate concepts from ejector mechanisms while selectively retaining and discarding traditional fan elements. ... Rather than relying solely on large rotating blades, we might consider smaller impellers driven by the existing motor, which would introduce directed air streams. ... Retaining a compact motor system that uses lightweight, efficient turbines is crucial, as it allows for consistent airflow without excessive bulk. ... By strategically designing air inlets and channels, we can naturally guide air currents, even without a high-velocity driving fluid. Prioritizing safety, all moving parts should be enclosed within the casing, utilizing protective materials to eliminate the risk of contact with blades. ... Utilize computational fluid dynamics (CFD) to refine blade shapes and angles, enhancing air movement and improving efficiency. ... Incorporate guide vanes that naturally direct airflow, minimizing reliance on high-velocity inputs. This approach ensures smoother air pressure control and efficient cooling without adding mechanical complexity. ...
@@
In contrast,the interpretation of the impact of the number of agents is not straightforward. On one hand, the effect is generally positive for models without an internal state. First, the average performance improved significantly when the number of agents increased from one (Model A) to multiple (Models C and F). Second, when comparing the three-agent model (Model C) to the five-agent model (Model F), increasing the number of agents significantly reduced the variance, although the average scores remained unchanged. The larger variance in scores with three agents may partly stem from the stricter majority requirement for consensus, where the opinion of a single agent can heavily influence the results.
@@
On the other hand, when the agents have an internal state, the single agent model (Model B) outperformed almost all multiple-agent models both with and without an internal state (Models C, D, E, F, and G) except for Model H. It also recorded the smallest variance across all models. This highlights the strong capability of the internal state module in deriving comprehensive and coherent ideas, even when acting alone. However, its superior performance compared to some multi-agent models with internal states may also be attributed to the stricter majority requirement. In particular, the larger variance observed in the scores with three agents suggests this.
@@
The highest average score among all models was recorded by the case with five heterogeneous agents with an internal state (Model H). Increasing the number of agents from 3 to 5 slightly relaxed the consensus requirement, allowing a broader range of opinions to emerge from more diverse agents, which were then refined and shared through collective reasoning. The positive effects of the number and diversity of agents are consistent with the results of some prior studies on LLM-MAAS ([du2023improvingfactualityreasoninglanguage, chan2023chatevalbetterllmbasedevaluators]). Below is an example of part of agents’ statement in the multiple-agent model with an internal state (Model H).
@@
Agent_1: ... First, retaining the electric motor is crucial, but shifting the focus from blades to a ducted design with specialized nozzles could minimize safety risks and maintenance. ... This means adopting a circular or linear duct system to channel airflow, creating efficient fluid movement pathways. The key is to enhance the fan’s ability to draw in and expel air smoothly using controlled high-speed airflow, perhaps with adjustable nozzles or diffusers that modify air velocity and direction. ...
Agent_2: ... However, traditional metal or plastic grids could be phased out if the airflow generation can be enclosed entirely within the fan’s structure.
Agent_3  ... The most critical adaptation would be incorporating the ejector’s pressure-driven flow to create a bladeless fan. This could be achieved using a small compressor or blower to generate a high-pressure air source, channeling it through a nozzle to produce a coherent airflow. ...
Agent_5: ... Retaining the motor to power a compressor or blower generating a directed airflow stream aligns with fluid dynamics principles, enhancing efficiency. ...
Agent_1: ... Noise minimization can be achieved by including sound-dampening materials and CFD-aided design to manage high-speed airflow noise impacts. ...
@@
The agents’ dialogue above proposed several ideas, including eliminating the existing grid guard and conducting airflow generation within an enclosed environment inside the fan, adopting a ducted design to improve airflow efficiency, utilizing a motor-driven blower to generate high-speed driving fluid, and performing CFD simulations to reduce noise. These proposals demonstrate that the key ideas underlying the Dyson Air Multiplier have been successfully replicated.
Model Min Max Average Standard deviation
A: 1 no-internal agent 1 5 2.4 1.58
B: 1 with-internal agent 4 7 5.1 0.99
C: 3 no-internal agents 1 7 3.7 2.45
D: 3 with-internal homogeneous agents 2 7 4.4 2.07
E: 3 with-internal heterogeneous agents 2 7 4.0 2.05
F: 5 no-internal agents 1 6 3.7 1.83
G: 5 with-internal homogeneous agents 2 6 4.7 1.16
H: 5 with-internal heterogeneous agents 3 7 5.3 1.16
TABLE III: Experimental results
IV Conclusion

### v3

@@ 3 Is retaining blades proposed, and is enclosing them inside the main body suggested? (1 point)
4 If no direct mention of a motor is made, is retaining blades proposed, or is the device proposed in criterion 1 conventionally paired with a motor? (1 point) Additionally, is the use of household electricity as the motor’s power source mentioned? (1 point)
5 Is retaining a motor proposed, and is it identified as playing a role in generating driving fluid or negative pressure? (1 point)
@@ 7 Is the introduction of a nozzle structure proposed? (1 point)
8 Other than the nozzle, are any proposals made to optimize internal channels for the efficiency or noise reduction of the driving fluid? (Up to 2 points depending on the number and specificity of the proposals)
@@ 9 Are relevant principles of fluid dynamics explicitly explained? (1 point)
To evaluate the performance of the proposed framework, the results are compared across eight models shown in Table 2, differentiated based on three perspectives: (1) the number of agents, (2) the presence of internal states, and (3) the heterogeneity of intrinsic motivations. The first one examines the impact of increasing the number of agents. When multiple agents are involved, the evaluation is conducted based on the majority opinions. The second one assesses the functionality of the internal state module. In models without internal states, agents are required to make statements based solely on their initially assigned roles and personalities, as well as the history of past conversations, without undergoing the process of reflection. The third one evaluates the effect of heterogeneity among agents. In homogeneous models (Models D and G), all agents equally consider the three intrinsic motivations (novelty, importance, and consensus). In heterogeneous models (Models E and H), agents are configured with different emphases: some prioritize novelty while de-emphasizing consensus, others prioritize consensus while de-emphasizing novelty, and some consider all three motivations equally. The experiment is conducted 10 times for each model using a common set of 10 seeds. The ratings are done by carefully examining the generated solutions through human assessment.
Table 2: Eight models for experiment
Model # of agents Internal state intrinsic motivation
Model A 1 No n.a.
Model B 1 Yes n.a.
Model C 3 No n.a.
Model D 3 Yes homogeneous
Model E 3 Yes heterogeneous
Model F 5 No n.a.
Model G 5 Yes homogeneous
Model H 5 Yes heterogeneous
3.3 EXPERIMENT 2
To examine the applicability of the framework, the second experiment applies it to ten different source domains: (1) safety sensors, (2) cooling mats, (3) shock-absorbing materials, (4) air coolers, (5) aerogenerators, (6) mist spray cooling systems, (7) small hydropower systems, (8) noise-canceling technology, (9) airbags, and (10) vacuum cleaners. These domains are selected because they share certain functions with household fans, such as cooling effects and energy utilization through rotational motion, or because they relate to key issues in household fans, such as injury prevention, noise reduction, and energy efficiency. The list of the ten domains and the reasons for selection is given in Table 3. The experiment is conducted for two extreme models: the sigle-agent model without an internal state (Model A) and the five heterogeneous agent model with internal states (Model H).
Table 3: Ten source domains
Domain Reason for selection
Safety sensors Related to key issues in household fans (injury prevention)
Cooling mats Share functions with household fans (cooling effect)
Shock-absorbing materials Related to key issues in household fans (injury prevention)
Air coolers Share functions with household fans (cooling effect)
Aerogenerators Share functions with household fans (energy utilization through rotational dynamics)
Mist spray cooling systems Share functions with household fans (cooling effect)
Small hydropower systems Share functions with household fans (energy utilization through rotational dynamics)
Noise-canceling technology Related to key issues in household fans (noise reduction)
Airbags Related to key issues in household fans (injury prevention)
Vacuum cleaners Share functions with household fans (energy utilization through rotational dynamics)
@@ 3.3 EXPERIMENT 2
In both experiments, the language model used to power the framework is OpenAI’s GPT-4o. To ensure uniform experimental conditions, word limits are imposed on the agents’ statements and reflections in each round. In this preliminary experiment, a flat organizational structure was adopted, where all agents could communicate with all others. Other structures, such as hierarchical models, are left for future research.
3.4 RESULTS
@@
Table 4 summarizes the results of Experiment 1. In all cases with 1, 3, and 5 agents, the models with internal states outperformed those without on average. This indicates that, through refinement by internal states, more coherent and comprehensive ideas are constructed and shared among agents. For example, in the single-agent model without an internal state (Model A), the proposal of retaining blades or incorporating an appropriate blade-driven device as a mechanism to generate a driving fluid (Criteria 1 and 2 in Table 1) was made only 3 times out of 10. In contrast, in the single-agent model with an internal state (Model B), such mechanisms were proposed 9 times out of 10. This indicates, in the latter model, the concept of external driving fluid was successfully reframed within the context of household fans as the intake of surrounding air by blades. Furthermore, in the model without an internal state, aside from the introduction of nozzle structures, the proposal for optimizing internal channels to improve the efficiency and reduce the noise of the driving fluid (Criterion 8) was made only once out of 10 times. In contrast, in the model with an internal state, such proposals were made 7 times out of 10, ranging from the redesign of the fan housing to the introduction of guide vanes and the use of Computational Fluid Dynamics (CFD) simulations. Below is an example of part of an agent’s statement in the single-agent model with an internal state (Model B).
@@
Agent_1: Rather than relying solely on large rotating blades, we might consider smaller impellers driven by the existing motor, which would introduce directed air streams. … Retaining a compact motor system that uses lightweight, efficient turbines is crucial, as it allows for consistent airflow without excessive bulk. … By strategically designing air inlets and channels, we can naturally guide air currents, even without a high-velocity driving fluid. Prioritizing safety, all moving parts should be enclosed within the casing, utilizing protective materials to eliminate the risk of contact with blades. … Utilize computational fluid dynamics (CFD) to refine blade shapes and angles, enhancing air movement and improving efficiency. … Incorporate guide vanes that naturally direct airflow, minimizing reliance on high-velocity inputs. This approach ensures smoother air pressure control and efficient cooling without adding mechanical complexity.
@@
In contrast, the impact of the number of agents is not straightforward. On one hand, the effect is generally positive for models without internal states. First, the average performance improved significantly when the number of agents increased from one (Model A) to multiple (Models C and F). Second, when comparing the three-agent model (Model C) to the five-agent model (Model F), increasing the number of agents significantly reduced the variance in scores, although the average scores remained unchanged. The larger variance in scores with three agents may partly stem from the stricter majority requirement for consensus, where the opinion of a single agent can heavily influence the results.
@@
On the other hand, when the agents have an internal state, the single-agent model (Model B) outperformed almost all multi-agent models both with and without internal states (Models C, D, E, F, and G) except for Model H. It also recorded the smallest variance across all models. This highlights the strong capability of the internal state module in deriving comprehensive and coherent ideas, even when acting alone. However, its superior performance compared to some multi-agent models with internal states may also be attributed to the stricter majority requirement. The larger variance observed in the three-agent models suggests this.
@@
The highest average score among all models was recorded by the case with five heterogeneous agents with internal states (Model H). Increasing the number of agents from 3 to 5 slightly relaxed the consensus requirement, allowing a broader range of opinions to emerge from more diverse agents, which were then refined and shared through collective reasoning. The positive effects of the number and diversity of agents are consistent with the results of some prior studies on LLM-MAAS ([7, 8, 16]). Below is an example of part of agents’ statements in the five heterogeneous agent model with internal states (Model H).
@@
Agent_1: … First, retaining the electric motor is crucial, but shifting the focus from blades to a ducted design with specialized nozzles could minimize safety risks and maintenance. … This means adopting a circular or linear duct system to channel airflow, creating efficient fluid movement pathways. The key is to enhance the fan’s ability to draw in and expel air smoothly using controlled high-speed airflow, perhaps with adjustable nozzles or diffusers that modify air velocity and direction. … Agent_2: … However, traditional metal or plastic grids could be phased out if the airflow generation can be enclosed entirely within the fan’s structure. Agent_3: … The most critical adaptation would be incorporating the ejector’s pressure-driven flow to create a bladeless fan. This could be achieved using a small compressor or blower to generate a high-pressure air source, channeling it through a nozzle to produce a coherent airflow. … Agent_5: … Retaining the motor to power a compressor or blower generating a directed airflow stream aligns with fluid dynamics principles, enhancing efficiency. … Agent_1: … Noise minimization can be achieved by including sound-dampening materials and CFD-aided design to manage high-speed airflow noise impacts. …
@@
The agents’ dialogue above proposed several ideas, including eliminating the existing grid guard and conducting airflow generation within an enclosed environment inside the fan, adopting a ducted design to improve airflow efficiency, utilizing a motor-driven blower to generate high-speed driving fluid, and performing CFD simulations to reduce noise. These proposals demonstrate that the key ideas underlying Dyson Air Multiplier have been successfully replicated.
Table 4: Results of Experiment 1
Model Min Max Average score Standard deviation
Model A: 1 agent without an internal state 1 5 2.4 1.58
Model B: 1 agent with an internal state 4 7 5.1 0.99
Model C: 3 homogeneous agents without internal states 1 7 3.7 2.45
Model D: 3 homogeneous agents with internal states 2 7 4.4 2.07
Model E: 3 heterogeneous agents with internal states 2 7 4.0 2.05
Model F: 5 homogeneous agents without internal states 1 6 3.7 1.83
Model G: 5 homogeneous agents with internal states 2 6 4.7 1.16
Model H: 5 heterogeneous agents with internal states 3 7 5.3 1.16
Table 5 summarizes the results of Experiment 2. The most significant difference between the two models is the levels of technological coherence. In the single-agent model without an internal state (Model A), several proposals for some source domains are highly inconsistent or misaligned with the intended applications. For instance, in the cases of cooling mats and shock-absorbing materials, the model suggests that incorporating these materials into the guard can reduce the risk of injury. Similarly, in the airbag case, it proposes a retractable blade system to store the blades when not in use, arguing that this could reduce the risk of accidental contact and potentially eliminate the need for a protective guard. Furthermore, in the safety sensors case, both models propose the introduction of an automatic stopping mechanism. However, while the five-agent model with internal states (Model H) suggests retaining the guard as a safety backup, the model without an internal state (Model A) considers the possibility of removing the guard entirely. In addition to the technological coherence, another noteworthy aspect of the five-agent model with internal states (Model H) is the diversity of proposals across multiple source domains. These include hybrid manual-automatic systems and the adoption of lightweight, durable materials, demonstrating a broader range of innovative solutions.
In sum, the second experiment demonstrates that the five heterogeneous agent model with internal states outperforms the single-agent model without an internal state both in the levels of technological coherence and diversity of proposals.
Table 5: Results of Experiment 2
Domain Common proposals Major differences
Model A Model H
Safety sensors Adaptive control systems that allow for autonomous real-time adjustments by smart sensors. Automatic stopping by proximity sensors. The conventional guard design can be discarded or modified. Maintaining a minimal guard design as a safety backup.
Cooling mats Passive cooling materials within fan structures to enhance cooling efficiency and reduce noise. The conventional guard design can be discarded or modified. Maintaining a minimal guard design as a safety backup.
Shock-absorbing materials Shock-absorbing materials in fan components to reduce noise. Adopting hock-absorbing materials for blade guards can mitigate the risk of injury. Adopting passive materials to absorb vibrations can enhance the fan’s energy efficiency.
Air coolers Incorporating evaporative cooling components. A bladeless design could be explored. Retaining core air movement features like rotating blades. A detachable, modular evaporative units with lightweight, eco-friendly materials.
Aerogenerators Sensors and automated systems to optimize airflow dynamically. Adjustable, aerodynamically efficient blade designs.
Mist spray cooling systems Hybrid systems integrating adjustable misting functions with traditional fan mechanisms. Modular and detachable misting components. Smart sensors to automatically adjust mist output. Moisture-resistant materials and waterproof seals.
Small hydropower systems Aadaptive control systems that allow for autonomous real-time adjustments by smart sensors. Advanced blade aerodynamics from hydropower turbines to design fan blades. Hybrid control mechanisms combining manual and automatic modes.
Noise-canceling technology Adaptive control systems that allow for autonomous real-time adjustments by smart sensors. Modifying blade materials to incorporate sound-dampening qualities. Hybrid control mechanisms combining manual and automatic modes. A modular design approach. Leverage machine learning to predict and adapt fan’s settings for optimal comfort. Motors utilizing principles from noise-canceling technology to reduce operational noise.
Airbags A retractable blade system that allows blades to retract into a housing unit when not in use, minimizing the risk of accidental contact. Automatic stopping mechanism by proximity sensors. The bulky guards could potentially be discarded in favor of these advanced safety mechanisms. Adaptive control systems that allow for autonomous real-time adjustments by smart sensors. Hybrid control mechanisms combining manual and automatic modes. Adopting lightweight, durable materials.
Vacuum cleaners Transitioning to enclosed or bladeless designs by reimagining the internal airflow pathway. Incorporating adjustable components that can customize airflow (e.g., directional louvers or nozzles). Adopting filtration technologies. Advanced brushless motor systems.
4 RELATED WORKS
4.1 INNOVATION and GENERATIVE AI
Numerous studies have examined the impact of generative AI on corporate innovation, particularly in the field of management science ([2, 4, 5, 6, 13, 15, 22, 23, 29, 30, 31, 57, 69, 71, 75, 74, 81, 82, 94]). These studies include analyses of the roles of generative AI in innovation management ([15, 57, 71]), experiments using generative AI in human’s innovation processes ([2, 4, 6, 23, 30, 81, 94]), the impact on human decision-making ([29]), the acceptance of generative AI ([13]), and the classification of innovation strategies ([31]). Additionally, numerous studies have explored the use of generative AI in patent analysis and technology assessment, such as patent mapping and technology opportunity evaluation ([11, 62]) as well as similarity analysis of human ideas ([40, 85]). However, none of these studies propose a system of generative agents itself that can be directly applied to innovation processes performed only by AI. Reference [24] evaluates the idea generation capabilities of LLMs using zero-shot and few-shot prompting; however, it does not incorporate interactions among multiple agents or agents with internal states. Reference [64] simulates human cultural evolution by conducting experiments in which groups of LLMs play Little Alchemy 2. However, the primary focus is on knowledge graph exploration rather than technological or corporate innovation. Additionally, it does not involve the design of agents with internal states.
Apart from technological and corporate innovation, numerous studies have examined the creative capabilities of generative AI, particularly in areas such as image generation. For instance, [14] attempts to analyze the influence of cultural biases on human innovation process using image generation AI. However, it does not utilize LLM agents with internal states. Additionally, [38] simulates the systems model of creativity as dynamic interactions among three entities: visual agent artist that creates paintings, field critic that evaluates them, and domain that indicates evaluation trends. Although [38] does not focus on technological or corporate innovation, the roles of the external field critic are similar to the introspection function of this study.
4.2 AGENT ARCHITECTURE
The agent architecture in this study is inspired by several studies on LLM-based agents, including multi-agent systems (for review articles, see [10, 27, 36, 47, 79, 89, 84, 99, 101]). First, many studies have incorporated a memory module ([12, 21, 34, 41, 46, 49, 52, 54, 60, 65, 66, 67, 68, 70, 72, 73, 83, 87, 93, 97, 96, 100, 102, 104]). For LLM-based agents, a memory system is crucial not only for mimicking human cognitive processes but also for facilitating self-evolution through experience accumulation or knowledge extraction ([99]). In this study, the memory module plays a key role in ensuring context-awareness and consistency in dialogue, as well as promoting the self-evolution of thought. In particular, as in studies such as [12, 46, 67, 93, 96], the memory module of this study stores reflections in addition to observations (documents and conversations), aiming to enhance idea refinement and knowledge extraction through the recurrent process of thought and dialogue. On the other hand, since the proposed framework is mainly designed for short-term discussions on focused topics, it does not introduce memory operations such as selection based on recency, importance, relevance, or similarity ([34, 41, 46, 65, 67, 96]), embedding vectorization ([34, 41, 49, 52, 66, 96, 102]), or distinction between short-term and long-term memory ([41, 68, 73, 93]).
Apart from memory modules, numerous studies have focused on designing the internal processes of agents that serve as the basis for decision-making ([12, 41, 46, 45, 44, 49, 55, 67, 73, 92, 97, 98]). Among these, this study is characterized by three key features. First, the internal state module in this study adopts a recursive structure to update agents’ current thoughts iteratively. Due to this structure, even without explicit memory operations such as memory selection, critical insights from past reflections or conversations are refined into higher-level inferences, strongly influencing agents’ future reflections and statements. Similar structures are employed in studies such as [41, 44]. Second, this study incorporates an introspection process within the internal state module, functioning as an inner critic. Other studies that have integrated inner critic mechanisms include the self-verification module in [83], reflection mechanism in [46], Refine step in [55], and evaluator’s role in [73]. Additionally, some studies have positioned the critic function externally to the agent, such as the Critic-In-The-Loop mechanism in [43] and Field critic in [38]. Third, this study adopts a structure in which diverse intrinsic motivations are reflected in the idea generation process. This ensures that the dialogue process balances both innovation and cooperation, which are essential components of real-world innovation processes. To my knowledge, no existing study has adopted such a structure for idea generation.
5 CONCLUSIONS